If the delivery sheet shows that you should have 6 crates of bread, each containing 12 loaves, how many loaves of bread short are you if only 3 crates are delivered?
Answer by math_helper(2461)


There are 12 loaves in one crate.
There are 24 (=2*12) loaves in two crates.
There are 36 (=3*12) loaves in three crates.
:
There are 72 (=6*12) loaves in 6 crates (just 6 times the one crate scenario).
:
In general, if you have N crates, you have 12*N loaves.

If 3 crates are delivered, that means you are 6-3=3 crates short.
So you are 3*12 = 36 loaves short.

I hope this makes sense to you. The important part here is to recognize that the number of loaves is directly related to the number of crates.